OpenStudy (anonymous):

Solve the equation. 1.9x - 3 - 1.4x = 14

OpenStudy (anonymous):

x=34

OpenStudy (anonymous):

x=34

OpenStudy (anonymous):

0.5x = 17 x = 34

OpenStudy (anonymous):

1.9x-1.4x=14+3 0.5x=17 x=17/0.5 x=34